The Blue line is the key answer.
A 2.5-mm diameter copper wire carries a 33-A curre
Answer #1

here,

current , I = 33 A

dimaeter , d = 2.5 * 10^-3 m

radius , r = 1.25 * 10^-3 m

a)

the magnetic feild at the surface , B = u0 * I /(2 * pi * r)

B = 1.26 * 10^-6 * 33 /( 2 * pi * 1.25 * 10^-3)

B = 5.3 * 10^-3 T

b)

x = 0.5 * 10^-3 m

so , x' = r - x = 0.75 * 10^-3 m

using Ampere's circuital law

Iin = I* ( pi * x^2) /( pi * r^2)

Iin = 33 * ( 0.75 * 10^-3)^2 /( 1.25 * 10^-3)^2

Iin = 11.88 A

the magnetic feild inside the wire , B' = u0 * Iin /( 2 * pi * x)

B' = 1.26 * 10^-6 * 11.88 /( 2 * pi * 0.75 * 10^-3)

B' = 3.2 * 10^-3 T